OpenStudy (anonymous):

s=4(pie)rsqured ...use properties of equality to find the radius of a sphere with a surface area of 36 square meters

OpenStudy (anonymous):

Let's substitute 36 = 4(pi) r squared and divide both sides by 4(pie) 9/pi = r^2 What next?

OpenStudy (anonymous):

square root?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yes, can you plug sqrt (9/pi) into your calculator?

OpenStudy (anonymous):

so 3/the square root of pie?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Yes ...

OpenStudy (anonymous):

thank you!

OpenStudy (across):